Magnesium hydroxide, the active ingredient in milk of magnesia, neutralizes stomach acid, primarily HCl, according to the reaction: 

                                Mg(OH)2 + 2 HCl à 2 H2O + MgCl2        

How much HCl in grams can be neutralized by 11.4 g of Mg(OH)2? 

Molar masses, g/mol: Mg(OH)2 = 58.33; HCl = 36.46


1
Expert's answer
2021-03-18T02:58:05-0400

Moles of "Mg(OH)_2" in "11.4\ g=\frac{11.4}{58.33}=" "0.19544"

From the equation,

"Mg(OH)_2 + 2 HCl\to2 H_2O + MgCl_2"

So,for every mole of magnesium hydroxide ,two moles of "HCl" are required.

Hence moles of "HCl" required are "0.19544\times 2=0.39088"

Mass of "HCl" required "=0.39088\times 36.46=14.251\ g"